OpenStudy (anonymous):

An aircraft is flying horizontally at a constant height of 4,000 ft above a fixed observation point. At a certain instant, the speed of the aircraft is 300 mi/hr and the angle of elevation is 30 degrees. a) How fast is theta decreasing at this instant? b)How fast is the distance between the aircraft and the observation point changing at this instant? Express the result in unites of ft/s. Remember, 1 mile is 5280 ft.

OpenStudy (anonymous):

Related rates, right?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

\[ \tan \theta = \frac{4,000}{x} \]

OpenStudy (anonymous):

Differentiate both sides.

OpenStudy (anonymous):

Then you get this tanΘ=4000x^−1

OpenStudy (anonymous):

Yeah. That is our equation. Now just differentiate.

OpenStudy (anonymous):

With respect to \(t\).

OpenStudy (anonymous):

sec^2θ (dθ/dt)=−4000x^−2 (dx/dt)

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

I got dθ/dt=−11/400

OpenStudy (anonymous):

"the speed of the aircraft is 300 mi/hr and the angle of elevation is 30 degrees. " This means \(\theta=30^\circ\) and \(x'=300\).

OpenStudy (anonymous):

What did you get for \(x\)?

OpenStudy (anonymous):

x=4000√3

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

So is dθ/dt correct?

OpenStudy (anonymous):

Don't you change 300 mi into feet because the height is 4000 ft

OpenStudy (anonymous):

So that the units are the same

OpenStudy (anonymous):

Oh, I didn't notice.

OpenStudy (anonymous):

-11/400 in radians.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think you can do these problems.

OpenStudy (anonymous):

I do not know how to do the next part

OpenStudy (anonymous):

I'm not sure if I go the right answer

OpenStudy (anonymous):

*got

OpenStudy (anonymous):

\[ z^2=x^2+y^2 \]Differentiate.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Thanks! :)